OpenStudy (anonymous):

The volume of a cylinder is 289 cm2 and the height is 21 cm. Find the diameter. Round the answer to the nearest whole number. A. 2 cm B. 3 cm C. 6 cm D. 4 cm

OpenStudy (anonymous):

ok what is the formula for cylinder

OpenStudy (anonymous):

Idk this is all it showed me

OpenStudy (anonymous):

\[v=\pi*r^{2}*h\] can you take it from here?

OpenStudy (anonymous):

something is wrong with this we website today

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (anonymous):

\[289 = \pi*21*r^{2}\] solve for r, can you?

OpenStudy (anonymous):

you still not done with it? how can you keep r on one side of the equation?

OpenStudy (anonymous):

i just want the answer but it doesnt matter anymore i already got my assignment graded

OpenStudy (anonymous):

\[289 = \pi * 21 *r^{2}\] \[\frac{ 289 }{ \pi * 21 } = \frac{ \pi * 21 *r^{2} }{ \pi * 21 }\] \[\frac{ 289 }{ \pi * 21 } = r^{2} \] \[\sqrt{\frac{ 289 }{ \pi * 21 }} = \sqrt{ r^{2} }\] \[\sqrt{4.38055} = r\] \[2.093 = r\] \[d=2*r\] \[d=2*(2.093)\] \[d=4.18 = 4cm\]

OpenStudy (anonymous):

was it hard